# Physics-Driven Data Generation for Contact-Rich Manipulation via Trajectory Optimization

### v1

@@ Abstract
00footnotetext: *Equal Contribution. 1Computer Science and Artificial Intelligence Laboratory (CSAIL), Massachusetts Institute of Technology. 2Robotics and AI Institute. Correspondence to lujie@mit.edu.
@@ I Introduction
The emergence of foundation models has transformed fields such as natural language processing and computer vision, where models trained on massive, internet-scale datasets demonstrate remarkable generalization across diverse reasoning tasks [1, 2, 3, 4, 5]. Motivated by this success, the robotics community is currently pursuing foundation models for generalist robot policies capable of flexible and robust decision-making across a wide range of tasks [6, 7, 8], leading to significant industrial investments in large-scale robot learning [9]. However, the pursuit for generalist robot policies remains constrained by the limited availability of high-quality datasets, especially for contact-rich robotic manipulation. Existing datasets [7, 10, 11, 12] are orders of magnitude smaller than those used to train foundation models in other domains, such as Large Language Models (LLMs). The scarcity of diverse, high-fidelity manipulation data limits policy generalization across different embodiments, task contexts, and physical conditions.
To address data scarcity, robot learning researchers often rely on a spectrum of data sources varying in cost, quality, and transferability. The most informative data typically consists of high-quality demonstrations specific to the task, environment, and embodiment [7, 10], but such data is costly and time-consuming to collect, as it requires human teleoperation with specialized hardware. At the opposite end of the spectrum, there is a wealth of lower-quality data in the form of internet videos showing humans and robots performing manipulation tasks [13, 14, 15, 16]. However, the significant embodiment gap and limited action labeling make this data difficult to transfer effectively to robot policies. Simulation data offers a middle ground, providing the potential to generate large, diverse, and high-quality datasets at relatively low cost [17, 18, 19]. In practice, effective policy learning can be achieved by co-training on a mixture of data from different points along this spectrum, reducing data collection costs while improving generalization [20].
Figure 1: Physics-driven data generation overview. Leveraging trajectory optimization, our framework automatically generates thousands of dynamically feasible contact-rich trajectories across a range of embodiments and physical parameters from only 24 human demonstrations. The policy trained with imitation learning from the generated dataset is more robust and performant.
@@
In this work, we propose a data generation framework that leverages the strengths of both approaches: human demonstrations can provide global guidance, while trajectory optimization can locally refine these demonstrations to ensure dynamic feasibility. Starting with a small number of human demonstrations collected in a virtual reality (VR) environment, our method uses model-based trajectory optimization to generate large datasets of dynamically feasible, contact-rich trajectories in simulation. The demonstrations guide the planner through complex search spaces, while the planner ensures physical consistency and robustness across varying physical parameters and robot embodiments. Our pipeline, visualized in Fig. 1, enables efficient cross-embodiment data transfer, where demonstrations collected with one robot configuration can be adapted to another, and supports domain randomization for improved generalization and robustness. Additionally, it provides the potential to revive and adapt legacy datasets collected with different hardware or configurations, making old datasets valuable for new robot systems.
@@ IV-B Demonstration-Guided Trajectory Optimization
Here, $f$ is obtained by time-stepping the dynamics engine, $x_{\text{min}}/x_{\text{max}}$ ($u_{\text{min}}/u_{\text{max}}$) are the lower and upper bounds on the state (input), $Q_{t},R_{t}$ are the cost matrices for the state and input, respectively, and $Q_{T}$ is the cost matrix for the terminal state. To encourage precise tracking of the object trajectory, we assign higher weights to the entries of $Q_{t}$ which correspond to $x_{t}^{\text{object}}$. The detailed parameters can be found in Appendix IX-A.
@@
1 Input: Probability distribution $\rho$, augmentation number $N$, demo trajectory $x^{\text{demo}}_{0:T}$;
2 Output: $N$ dynamically consistent trajectories on target embodiments $\{(x_{0:T}^{\star},u_{0:T-1}^{\star})\}$;
@@ 5       Sample $\theta_{0:T}\sim\rho$;
Init. obj. trans. pert. (cm) [±$\pm 1.5$, ±$\pm 1.5$, 0] [±$\pm 5$, ±$\pm 5$, 0]
Init. obj. rot. pert. (rad) [0, 0, ±$\pm 0.3$] [0, 0, ±$\pm 0.3$]

### v2

@@ Abstract
00footnotetext: * Equal contribution. Correspondence to <<lujie@mit.edu>>.
@@ I Introduction
The emergence of foundation models has transformed fields such as natural language processing and computer vision, where models trained on massive, internet-scale datasets demonstrate remarkable generalization across diverse reasoning tasks [1, 72, 2]. Motivated by this success, the robotics community is currently pursuing foundation models for generalist robot policies capable of flexible and robust decision-making across a wide range of tasks [34, 50, 51], leading to significant industrial investments in large-scale robot learning [63]. However, the pursuit for generalist robot policies remains constrained by the limited availability of high-quality datasets, especially for contact-rich robotic manipulation. Existing datasets [50, 33, 74, 15] are orders of magnitude smaller than those used to train foundation models in other domains, such as Large Language Models (LLMs). The scarcity of diverse, high-fidelity manipulation data limits policy generalization across different embodiments, task contexts, and physical conditions.
To address data scarcity, robot learning researchers often rely on a spectrum of data sources varying in cost, quality, and transferability. The most informative data typically consists of high-quality demonstrations specific to the task, environment, and embodiment [50, 33], but such data is costly and time-consuming to collect, as it requires human teleoperation with specialized hardware. At the opposite end of the spectrum, there is a wealth of lower-quality data in the form of internet videos showing humans and robots performing manipulation tasks [14, 61, 45, 32]. However, the significant embodiment gap and limited action labeling make this data difficult to transfer effectively to robot policies. Simulation data offers a middle ground, providing the potential to generate large, diverse, and high-quality datasets at relatively low cost [80, 5, 30]. In practice, effective policy learning can be achieved by co-training on a mixture of data from different points along this spectrum, reducing data collection costs while improving generalization [76, 78].
@@
In this work, we propose a data generation framework that leverages the strengths of both approaches: human demonstrations can provide global guidance, while trajectory optimization can locally refine these demonstrations to ensure dynamic feasibility. Starting with a small number of human demonstrations collected in a virtual reality (VR) environment, our method uses model-based trajectory optimization to generate large datasets of dynamically feasible, contact-rich trajectories in simulation. The demonstrations guide the planner through complex search spaces, while the planner ensures physical consistency and robustness across varying physical parameters and robot embodiments. Our pipeline, visualized in Fig. LABEL:fig:banner, enables efficient cross-embodiment data transfer, where demonstrations collected with one robot configuration can be adapted to another, and supports domain randomization for improved generalization and robustness. Additionally, it provides the potential to revive and adapt legacy datasets collected with different hardware or configurations, making old datasets valuable for new robot systems.
@@ IV-B Demonstration-Guided Trajectory Optimization
Here, $u$ is the control input, $f$ is obtained by time-stepping the dynamics engine, $x_{\text{min}}/x_{\text{max}}$ ($u_{\text{min}}/u_{\text{max}}$) are the lower and upper bounds on the state (input), $Q_{t},R_{t}$ are the cost matrices for the state and input, respectively, and $Q_{T}$ is the cost matrix for the terminal state. To encourage precise tracking of the object trajectory, we assign higher weights to the entries of $Q_{t}$ which correspond to $x_{t}^{\text{object}}$. The detailed parameters can be found in Appendix X-A.
@@
1Input: Probability distribution $\rho$, augmentation number NN, demo trajectory $x^{\text{demo}}_{0:T}$;
2 Output: NN dynamically consistent trajectories on target embodiments $\{(x_{0:T}^{\star},u_{0:T-1}^{\star})\}$;
@@ 5   Sample $\theta_{0:T}\sim\rho$;
Init. obj. trans. pert. (cm) $\pm 1.5$, $\pm 1.5$, 0] $\pm 5$, $\pm 5$, 0]
Init. obj. rot. pert. (rad) [0, 0, $\pm 0.3$] [0, 0, $\pm 0.3$]
@@ V Trajectory Optimization Experiments
We compare our method to MimicGen [41], an automatic data generation pipeline that adapts demonstrations to diverse spatial configurations through geometric transformations and replay. To ensure a fair comparison, we collect 24 demonstrations directly with the floating Allegro hand in simulation—eliminating embodiment mismatch between the source demonstration and target task—and apply MimicGen-style replay under varying physical parameters (Table I). Pure replay achieves a success rate of $51\%$, while trajectory optimization increases this to $86\%$, and smoothes jitters and imperfections in human demonstrations. Notably, our tasks are highly contact-rich, requiring joint-level control with frequent contact switches. In contrast, the contact behaviors in MimicGen (e.g., pick-and-place, insertion) involve fewer contact switches. Moreover, our task involves whole-body manipulation, requiring the entire robot to interact with the object. Simply transforming the end-effector pose in an object-centric manner as in MimicGen disregards the contact between the rest of the robot and the object, and can easily result in loss of contact and task failure.
@@ VIII Conclusion
IX Acknowledgement
The authors would like to thank Huaijiang Zhu, Haonan Chen and Jiayuan Mao for valuable discussions and insightful feedback on the paper. This work was supported by Robotics and AI Institute Agmd Dtd 8/1/2023, Amazon PO 2D-15694085 and PO 2D-15694048.